can anyone help me work this one out?
-x+3y=-1, x-2y=2
: can anyone help me work this one out?
-x+3y=-1, x-2y=2

Answer by edjones(2415)

x-2y=2
x=2y+2
.
-x+3y=-1
-(2y+2)+3y=-1 Substitute the 1st equation for x.
-2y-2+3y=-1
y-2=-1
y=1
x=4
.
Ed